mykenzie:

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represents 60 miles? Part B How many miles is represented by 1/4 inch on the map?

papya:

do you have a ruler at hand

papya:

4 inches = 50 miles so you just have to find what = 10 miles

papya:

1inch = 15 miles so how much shorter would be 10 miles

mykenzie:

no i don't have a ruler

papya:

how bout a calculator

mykenzie:

yes

papya:

i was wrong about 1=15miles it would be half of 25 miles

papya:

which would be like 12.5

mykenzie:

it is 12.5

papya:

to find your 1/4 you would multiply .25 by 12.5

mykenzie:

it is 3.12

papya:

that would be your 1/4

mykenzie:

ok

mykenzie:

thank you

papya:

then you should be able to find your 60 miles now

snowflake0531:

\[\frac{ 2 }{ 25 } = \frac{ x }{ 60 }\] solve for x by cross multiplication \[25x=120\] \[x=4.8\]

snowflake0531:

Then for the 1/4 of an inch, \[\frac{ 2 }{ 25 } = \frac{ \frac{ 1 }{ 4 } }{ x }\] cross multiply \[6.25 = 2x\] simplifies to \[x=3.125\]

mykenzie:

ok thank you for your help

snowflake0531:

you're welcome~